OpenStudy (anonymous):

log 3 (x2+8)-log 3 4 = 3

OpenStudy (callisto):

\[log_3 (x^2+8)-log_3 4 = 3\]\[log_3 \frac{(x^2+8)}{4} = 3\]\[\frac{(x^2+8)}{4} = 3^3\]\[x^2+8 = 108\]Can you continue?

OpenStudy (anonymous):

Yes, thank you :) Answer is 10 or -10.

OpenStudy (callisto):

welcome!~